Solve the equation:
4/x+2=12/x

Answer by TimothyLamb(4379)

4/x + 2 = 12/x
4/x - 12/x = -2
-8/x = -2
x = -8/-2
x = 4
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php